# A real-time digital twin of azimuthal thermoacoustic instabilities

### v1

@@ 6 Conclusions
We develop a real-time digital twin of azimuthal thermoacoustics of a hydrogen-based combustor for uncovering the physical state from raw data, and predicting the nonlinear dynamics in time. We identify the four ingredients to design a robust real-time digital twin:
@@
    In a real-time context, we need to work with raw data as they come from sensors, therefore, we do not preprocess the data. The data we utilize is raw (the data contain both environmental and instrumental noise, and turbulent fluctuations) and sparse (the data are collected by four microphones).
@@
    An estimator of both the model bias, which originates from the modelling assumptions and approximation of a low-order model; and measurement bias, which originates from microphones recording the total pressure (which has a non-zero mean) instead of the acoustic pressure (which has zero mean).
@@
We propose a real-time data assimilation framework to infer the acoustic pressure, physical parameters, model biases and measurements biases simultaneously, which is the bias-regularized ensemble Kalman filter (r-EnKF). We find an analytical solution that minimizes the bias-regularized data assimilation cost function (28). We propose a reservoir computer (echo state network) and a training strategy to infer both the model bias a measurement bias, without making assumptions on their functional forms a priori. The real-time digital twin is applied to a laboratory hydrogen-based annular combustor for a variety of equivalence ratios. The data are treated as if they came from sensors on the fly, i.e., the pressure measurements are assimilated at a time step and then disregarded until the next pressure measurement becomes available. First, after assimilation of raw data (at 1707 Hz, i.e., approximately 0.6 data points per acoustic period), the model learns the correct physical state and optimal parameters. The real-time digital twin autonomously predicts azimuthal dynamics beyond the assimilation window, i.e., without seeing more data. This is in stark contrast with state-of-the-art methods based on the bias-unregularized methods, which do not perform well in the presence of model and measurement biases. Second, the digital twin uncovers the physical acoustic pressure from the raw data. Because the physical mechanisms are constrained in the low-order model, which originate from conservation laws, the digital twin acts as a physics-based filter, which removes aleatoric noise and turbulent fluctuations. Third, physically we find that azimuthal oscillations are governed by a time-varying parameter system, which generalizes existing models that have constant parameters, and capture only slow-varying variables. Fourth, we find that the key parameters that influence the dynamics are the linear growth rates, and the angular frequency. Further, the digital twin generalizes to equivalence ratios at which current modelling approaches are not accurate. This work opens new opportunities for real-time digital twinning of multi-physics problems. Current work is focused on exploring the role of sparsity in the data.
@@ Appendix A The ensemble Kalman filter
As mentioned in §1, the EnKF is a bias-unregularized method. The estate and parameter update does not consider either the model bias or the measurement bias.
@@ Appendix B Jacobian of the bias estimator
$\bm{\mathsf{J}}=\dfrac{\mathrm{d}\,{\bm{\mathit{b}}}^{\mathrm{f}}_{k+1}}{\mathrm{d}\,\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}}=\dfrac{\mathrm{d}\,{\bm{\mathit{b}}}^{\mathrm{f}}_{k+1}}{\mathrm{d}\,\bar{{\bm{\mathit{i}}}}_{k}^{\mathrm{a}}}\dfrac{\mathrm{d}\,\bar{{\bm{\mathit{i}}}}_{k}^{\mathrm{a}}}{\mathrm{d}\,\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}}=-\dfrac{\mathrm{d}\,{\bm{\mathit{b}}}^{\mathrm{f}}_{k+1}}{\mathrm{d}\,\bar{{\bm{\mathit{i}}}}_{k}}=-\bm{\mathsf{W}}_{\mathrm{out}}^{(1)}\left[\bm{\mathsf{T}}\odot\left(\sigma_{\mathrm{in}}\bm{\mathsf{W}}_{\mathrm{in}}^{(1)}\odot\bm{\mathsf{G}}\right)\right],$ (38)
@@
${\bm{\mathit{T}}}={\bm{\mathit{1}}}-\tanh^{2}\left(\sigma_{\mathrm{in}}\bm{\mathsf{W}}_{\mathrm{in}}^{(1)}\left({\bm{\mathit{b}}}_{k}\odot{\bm{\mathit{g}}}\right)+\sigma_{\mathrm{in}}\delta_{\mathrm{r}}\bm{\mathsf{W}}_{\mathrm{in}}^{(2)}+\rho\bm{\mathsf{W}}{\bm{\mathit{r}}}_{k}\right).$
@@ Appendix C Training the network
During training, the ESN is in open-loop configuration (Figure 3a). The inputs to the reservoir are the training dataset $\bm{\mathsf{U}}=\left[{\bm{\mathit{u}}}_{0}\,|\dots|\,{\bm{\mathit{u}}}_{N_{\mathrm{tr}}-1}\right]$, where each time component ${\bm{\mathit{u}}}_{k}=[{\bm{\mathit{b}}}_{\mathrm{u}}(t_{k});{{\bm{\mathit{i}}}}_{\mathrm{u}}(t_{k})]$ (the subscript ‘u’ indicates training data). Although we have information from the experimental data to train the network, the optimal parameters of the thermoacoustic system are unknown. Thus, we do not know a priori the model bias and measurement bias. Selecting an appropriate training dataset is essential to obtain a robust echo state network which can estimate the model bias and innovations. We create a set of $L$ guesses on the bias and innovations from a single realization of the experimental data. This means that the ESN is not trained with the ‘true’ bias.
The training data generation is summarized in Algorithm 1, and the procedure is as follows. First, we take measurements for a training time window $t_{\mathrm{train}}$ of acoustic pressure data $\bm{\mathsf{D}}_{\mathrm{u}}$, and we estimate $\bm{\mathsf{D}}^{\dagger}_{\mathrm{u}}$ by filtering the PSD of $\bm{\mathsf{D}}$ (see §2). Second, we draw $L$ sets of thermoacoustic parameters from uniform random distribution with lower and upper bounds ${\bm{\mathit{\alpha}}}^{0}(1-\sigma_{L})$ and ${\bm{\mathit{\alpha}}}^{0}(1+\sigma_{L})$, which are selected from an educated physical initial guess on the thermoacoustic model parameters. Then, we forecast the model (7) to statistically stationary conditions using the $L$ sets of parameters, and we take from each time series a sample of length $t_{\mathrm{train}}$, with this, we have the model estimates $\bm{\mathsf{Q}}_{\mathrm{u}}\in\mathbb{R}^{L\times N_{q}\times N_{\mathrm{train}}}$. Third, we correlate each $\bm{\mathsf{Q}}_{\mathrm{u},l}$ with the data by selecting the time lag $\varkappa\geq 0$ which minimizes their normalized root mean square error $(\mathrm{RMS})$ within the first 0.01 s (approximately 10 periods of the acoustic signal), i.e., for each $\bm{\mathsf{Q}}_{l}$, the time lag is $\varkappa_{l}=\{\varkappa\;s.t.\;\min{\mathrm{RMS}(\bm{\mathsf{Q}}_{\mathrm{u},l}(t-\varkappa),\bm{\mathsf{D}}(t))}\}$ with $t\in[0,0.01]$ s. Once the $L$ model estimates are aligned to the observations such that the RMS is minimized in the initial 10 periods of oscillation, we obtain the training model bias and innovations as
@@
If our initial guess in the parameters is well defined, the training bias dataset $\bm{\mathsf{B}}_{\mathrm{u}}$ have small norm. However, during assimilation the system can go through different states and parameter combinations, which may have a large norm model bias. Because the echo state network is only trained on the correlated signals, it may not be flexible to estimate the bias throughout DA Liang et al. (2023). Therefore, fourth, we apply data augmentation (Goodfellow et al., 2016) to increase the robustness of the ESN and improve the network adaptive in a real-time assimilation framework. We increase the training set by adding the bias and innovations resulting from mid-correlated signals to the training set. We select the mid-correlation time lag as the average between the best lag $\varkappa$ and the worst time lag. The total number of training time series in the proposed training method is $2L$. The training parameters used in this work are summarized in D.
@@
1:${\bm{\mathsf{D}}_{\mathrm{u}}}\leftarrow$ Get observations(observation time = $t_{\mathrm{train}}$)
@@
6:     $\bm{\mathsf{Q}}\leftarrow$ Forecast model(${\bm{\mathit{\alpha}}}_{l},t_{\mathrm{train}}$)
@@ Appendix D Simulations’ parameters
This appendix summarizes the parameters used for the data assimilation and for training the echo state network.
@@
Assimilation $\Delta t_{\mathrm{d}}$ $35\Delta t$ s $N_{q}$ 4 $\sigma_{\mathrm{d}}$ 0.1
$\Delta t$ ${51200}^{-1}$ s $m$ 20 $\omega$a $[1090,1095]\cdot 2\pi$ Hz
$\nu$a $[-10,20]$ Hz $c_{2}\beta$a $[10,50]$ Hz $\kappa$a [1, 2]$\cdot 10^{-4}$ Hz
$\Theta_{\beta}$a $[0.5,0.7]$ rad $\Theta_{\epsilon}$a $[0.4,0.6]$ rad $\epsilon$a [5, 8] $\cdot 10^{-3}$
@@
$N_{\mathrm{wash}}$ 10 $t_{\mathrm{train}}$ 0.167 s $\rho$b $[0.5,1.0]$
@@
aIndicates that the parameters of the ensemble members are initialized within the given range.
@@ Appendix E Inferred thermoacoustic parameters
Table 2 details the parameters shown in Figures 9-10. The reported parameters from the real-time digital twin (DT) are those with minimum RMS.

### v2

@@ 7 Conclusions
We develop a real-time digital twin of azimuthal thermoacoustic oscillations of a hydrogen-based combustor for uncovering the physical state from raw data, and predicting the nonlinear dynamics in time. We identify the four ingredients requited to design a robust real-time digital twin:
@@
    In a real-time context, we need to work with raw data as they come from sensors, therefore, we do not pre-process the data. The data we utilize is raw (the data contain both environmental and instrumental noise, and turbulent fluctuations) and sparse (the data are collected by four microphones).
@@
    An estimator of both the model bias, which originates from the modelling assumptions and approximation of a low-order model; and measurement shift, which originates from microphones recording the total pressure (which has a non-zero mean) instead of the acoustic pressure (which has zero mean).
@@
We propose a real-time data assimilation framework to infer the acoustic pressure, physical parameters, model biases and measurements biases simultaneously, which is the bias-regularized ensemble Kalman filter (r-EnKF). We find an analytical solution that minimizes the bias-regularized data assimilation cost function (24). We propose a reservoir computer (echo state network) and a training strategy to infer both the model bias a measurement shift, without making assumptions on their functional forms a priori. The real-time digital twin is applied to a laboratory hydrogen-based annular combustor for a variety of equivalence ratios. The data are treated as if they came from sensors on the fly, i.e., the pressure measurements are assimilated at a time step and then disregarded until the next pressure measurement becomes available. First, after assimilation of raw data (at 1707 Hz, i.e., approximately 0.6 data points per acoustic period), the model learns the correct physical state and optimal parameters. The real-time digital twin autonomously predicts azimuthal dynamics beyond the assimilation window, i.e., without seeing more data. This is in stark contrast with state-of-the-art methods based on the bias-unregularized methods, which do not perform well in the presence of model bias and measurement shift. Second, the digital twin uncovers the physical acoustic pressure from the raw data. Because the physical mechanisms are constrained in the low-order model, which originate from conservation laws, the digital twin acts as a physics-based filter, which removes aleatoric noise and turbulent fluctuations. Third, physically we find that azimuthal oscillations are governed by a time-varying parameter system, which generalizes existing models that have constant parameters, and capture only slow-varying variables. Fourth, we find that the key parameters that influence the dynamics are the linear growth rates, and the angular frequency. Further, the digital twin generalizes to equivalence ratios at which current modelling approaches are not accurate. Fifth, to generalize the digital twin framework, we train an ESN with data from three of the four available equivalence ratios, and test the assimilation performance in the unseen scenario. This unified ESN successfully estimates the model bias and measurement shift in unseen thermoacoustic dynamics with the regularized bias-aware data assimilation framework. This work opens new opportunities for real-time digital twinning of multi-physics problems. Current work is focused on exploring the role of sparsity in the data.
@@ Appendix A The ensemble Kalman filter
As mentioned in §1, the EnKF is a bias-unregularized method. The estate and parameter update does not consider either the model bias or the measurement shift.
@@ Appendix B Jacobian of the bias estimator
$\bm{\mathsf{J}}=\dfrac{\mathrm{d}\,{\bm{\mathit{b}}}_{k+1}}{\mathrm{d}\,\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}}=\dfrac{\mathrm{d}\,{\bm{\mathit{b}}}_{k+1}}{\mathrm{d}\,\bar{{\bm{\mathit{i}}}}_{k}^{\mathrm{a}}}\dfrac{\mathrm{d}\,\bar{{\bm{\mathit{i}}}}_{k}^{\mathrm{a}}}{\mathrm{d}\,\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}}=-\dfrac{\mathrm{d}\,{\bm{\mathit{b}}}_{k+1}}{\mathrm{d}\,\bar{{\bm{\mathit{i}}}}_{k}}=-\bm{\mathsf{W}}_{\mathrm{out}}^{(1)}\left[\bm{\mathsf{T}}\odot\left(\sigma_{\mathrm{in}}\bm{\mathsf{W}}_{\mathrm{in}}^{(1)}\odot\bm{\mathsf{G}}\right)\right],$ (34)
@@
${\bm{\mathit{T}}}={\bm{\mathit{1}}}-\tanh^{2}\left(\sigma_{\mathrm{in}}\bm{\mathsf{W}}_{\mathrm{in}}^{(1)}\left({\bm{\mathit{b}}}_{k}\odot{\bm{\mathit{g}}}\right)+\sigma_{\mathrm{in}}\delta_{r}\bm{\mathsf{W}}_{\mathrm{in}}^{(2)}+\rho\bm{\mathsf{W}}{\bm{\mathit{r}}}_{k}\right).$
@@ Appendix C Training the network
During training, the ESN is in open-loop configuration (Figure 4a). The inputs to the reservoir are the training dataset $\bm{\mathsf{U}}=\left[{\bm{\mathit{u}}}_{0}\,|\dots|\,{\bm{\mathit{u}}}_{N_{\mathrm{tr}}-1}\right]$, where each time component ${\bm{\mathit{u}}}_{k}=[{\bm{\mathit{b}}}_{\mathrm{u}}(t_{k});{{\bm{\mathit{i}}}}_{\mathrm{u}}(t_{k})]$ (the subscript ‘u’ indicates training data). Although we have information from the experimental data to train the network, the optimal parameters of the thermoacoustic system are unknown. Thus, we do not know a priori the model bias and measurement shift. Selecting an appropriate training dataset is essential to obtain a robust echo state network which can estimate the model bias and innovations. We create a set of $L$ guesses on the bias and innovations from a single realization of the experimental data. This means that the ESN is not trained with the ‘true’ bias.
The training data generation is summarized in Algorithm 1, and the procedure is as follows. First, we take measurements for a training time window $t_{\mathrm{tr}}$ of acoustic pressure data $\bm{\mathsf{D}}_{\mathrm{u}}$, and we estimate $\bm{\mathsf{D}}^{\dagger}_{\mathrm{u}}$ by applying a band-pass filter to the raw data $\bm{\mathsf{D}}$ (see §2). Second, we draw $L$ sets of thermoacoustic parameters from uniform random distribution with lower and upper bounds ${\bm{\mathit{\alpha}}}^{0}(1-\sigma_{L})$ and ${\bm{\mathit{\alpha}}}^{0}(1+\sigma_{L})$, which are selected from an educated physical initial guess on the thermoacoustic model parameters based on previous works (Indlekofer et al., 2021; Faure-Beaulieu et al., 2021b). The ranges of the uniform distributions are reported in D. Then, we forecast the model (7) to statistically stationary conditions using the $L$ sets of parameters, and we take from each time series a sample of length $t_{\mathrm{tr}}$, with this, we have the model estimates $\bm{\mathsf{Q}}_{\mathrm{u}}\in\mathbb{R}^{L\times N_{q}\times N_{\mathrm{tr}}}$. Third, we correlate each $\bm{\mathsf{Q}}_{\mathrm{u},l}$ with the data by selecting the time lag $\varkappa\geq 0$ which minimizes their normalized root mean square error $(\mathrm{RMS})$ within the first 0.01 s (approximately 10 periods of the acoustic signal), i.e., for each $\bm{\mathsf{Q}}_{l}$, the time lag is $\varkappa_{l}=\{\varkappa\;s.t.\;\min{\mathrm{RMS}(\bm{\mathsf{Q}}_{\mathrm{u},l}(t-\varkappa),\bm{\mathsf{D}}(t))}\}$ with $t\in[0,0.01]$ s. Once the $L$ model estimates are aligned to the observations such that the RMS is minimized in the initial 10 periods of oscillation, we obtain the training model bias and innovations as
@@
If our initial guess in the parameters is well defined, the training bias dataset $\bm{\mathsf{B}}_{\mathrm{u}}$ have small norm. However, during assimilation the system can go through different states and parameter combinations, which may have a large norm model bias. Because the echo state network is only trained on the correlated signals, it may not be flexible to estimate the bias throughout the data assimilation process Liang et al. (2023). Therefore, fourth, we apply data augmentation (Goodfellow et al., 2016) to increase the robustness of the ESN and improve the network adaptive in a real-time assimilation framework. We increase the training set by adding the bias and innovations resulting from mid-correlated signals to the training set. We select the mid-correlation time lag as the average between the best lag $\varkappa$ and the worst time lag. The total number of training time series in the proposed training method is $2L$. The training parameters used in this work are summarized in D.
@@
1:${\bm{\mathsf{D}}_{\mathrm{u}}}\leftarrow$ Get observations(observation time = $t_{\mathrm{tr}}$)
@@
6:     $\bm{\mathsf{Q}}\leftarrow$ Forecast model(${\bm{\mathit{\alpha}}}_{l},t_{\mathrm{tr}}$)
@@ Appendix D Simulations’ parameters
This appendix summarizes the parameters used for the data assimilation and for training the echo state network. The model parameters’ ranges listed in Tab. 2 are used to initialize the ensemble in the data assimilation algorithm, and to create the $L$-initial guesses for training the ESN (see C).
@@
Assimilation $\Delta t_{d}$ $35\Delta t$ s $N_{q}$ 4 $\epsilon_{d}$ 0.1
$\Delta t$ ${51200}^{-1}$ s $m$ 20 Inflation 1.001
Model $\nu$a $[-10,30]$ Hz $c_{2}\beta$a $[10,50]$ Hz $\omega/(2\pi)$a $[1090,1095]$ Hz
$\Theta_{\epsilon}$a $[0.4,0.6]$ rad $\kappa\cdot 10^{4}$a $[1,2]$ Hz
$\Theta_{\beta}$a $[0.5,0.7]$ rad $\epsilon\cdot 10^{3}$a $[5,8]$
@@
$N_{\mathrm{wash}}$ 10 $t_{\mathrm{tr}}$ 0.167 s $\rho$b $[0.5,1.0]$
@@
aIndicates that the parameters are initialized within the given range.
@@ Appendix E Inferred thermoacoustic parameters
Table 3 details the parameters shown in Figures 10-11. The reported parameters from the real-time digital twin (DT) are those with minimum RMS.